OpenStudy (abbycross167):

Can somebody please just help me? I really need help.... Find the slope-intercept form of the equation that goes through these two points. (1,20) and (1,4.5)

pooja195 (pooja195):

Start by finding the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{4.5-20}{1-1}=slope\] Find the slope

OpenStudy (abbycross167):

15.5/0?

OpenStudy (abbycross167):

@pooja195

pooja195 (pooja195):

Wow i didn't realize what the end would be o-o infinty can't be a slope.... @sleepyjess ? @rvc

OpenStudy (abbycross167):

this is hard :/

pooja195 (pooja195):

x values are the same so its a vertical line

OpenStudy (abbycross167):

Yes ma'am / sir but how would I find the slope-intercept?

jigglypuff314 (jigglypuff314):

I don't believe you can get the slope intercept form if the slope is dividing by 0 your equation would be \(x = (number)\) to know what this number is, just look at the x values of the points they gave what are the x values of the points? :)

OpenStudy (abbycross167):

are they 1?

jigglypuff314 (jigglypuff314):

that's correct ^_^ so your answer should be x = 1

OpenStudy (abbycross167):

what? really? that's it?

jigglypuff314 (jigglypuff314):

yep but only because when you tried to find the slope - it was dividing by zero :)

OpenStudy (abbycross167):

Find the slope-intercept form of the equation that goes through these two points. (1,20) and (1,4.5) my Answer~ x=1

OpenStudy (abbycross167):

does that sound like the right answer?

jigglypuff314 (jigglypuff314):

nope because it is asking for slope-intercept form ^_^ but with the points given, that's the only answer that works

OpenStudy (baru):

yes x=1 is correct

OpenStudy (baru):

oh..yea, technically, it isn't the slope intercept form,

OpenStudy (abbycross167):

So that's what I need to put as the answer? x=1?

OpenStudy (abbycross167):

@baru and @jigglypuff314

TheSmartOne (thesmartone):

Question makes no sense, ask your teacher.

OpenStudy (baru):

Yes, put x=1 as ur answer If possible, add a note saying that we cannot put this line in slope intercept form, because its a vertical line. Vertical lines do not have a y intercept and have a slope of infinity.